Find the intercepts and then use them to graph the equation
1.1x-1.9y=4.18
I need to plot the points

Answer by Alan3354(69443)

Find the intercepts and then use them to graph the equation
1.1x-1.9y=4.18
-------------
To find the y-intercept, set x=0
-1.9y = 4.18
y = -2.2, so the intercept is (0,-2.2)
-----------------
For the x-intercept:
1.1x = 4.18
x = 3.8, so the x-intercept is (3.8,0)
Draw a straight line thru the points for the graph.